OpenStudy (anonymous):

Pls help quick. I've tried multiple different ways to get the answer and I never come up with the right one. Fan and Medal! Marylou is blowing up a beach ball. The package says that the ball has a diameter of 18 inches. What is the volume of the beach ball when it is completely full? Use 3.14 for . A. 339.12 cubic inches B. 3,052.08 cubic inches C. 24,416.6400000000 cubic inches D. 1,017.36 cubic inches

OpenStudy (anonymous):

lol, so c you say? Thanks. Lemme check the answer and get right back to you.

OpenStudy (anonymous):

Oops, turns out it was B. Thanks for trying tho! You still have a medal and new fan!

OpenStudy (anonymous):

Yup, it gives me an explanation. Here: Use the formula below to find the volume of the beach ball. Remember that the radius is one-half of the diameter. V=3/4 pi r^3 =4/3(3.14)(9 in)^3 =3,052.08 Therefore, the beach ball has a volume of 3,052.08 cubic inches when it is completely full.

OpenStudy (heyitslizzy13):

Just wanted to make sure you understood :) good job.